OpenStudy (anonymous):

Part A: Solve F = 9/5= C + 32 for C. Part B: Determine the value of C when F = 10 °F. Part C: Solve -np - 70 > 40 for n. Show your work

OpenStudy (anonymous):

Part A:9/5c+32=F -32 to both sides 9/5c=F-32 this is as far as I got in part A I don't know what I should do next

OpenStudy (anonymous):

to make 9/5c into just c (which is essentially 1c ) multiply each side by its reciprocal.....5/9

OpenStudy (anonymous):

So C=F-32/9/5